ore: 1097/3100 Resources Give up? Hint Consider the accompanying supply and demand graph. What is the value of consumer surplus? Supply What is the value of producer surplus? 5, 4.5) Demand What is the value of total (also called social or economic) surplus? Quantity
Answer #1

1). Consumer Surplus is the surplus which consumer gets when the market price of a good is less than the price consumer is willing to pay for that good.

Consumer Surplus = 1/2 * (Difference between the max price he is willing to pay and market price * Quantity)

Max price = 9, Market price = 4.5 , Quantity = 5

= 1/2 * ( 9 - 4.5 ) * 5

Value of consumer surplus = 11.25

2). Producer Surplus is the surplus which producer gets when the market price is higher than the price at which he is willing to sell his good.

Producer Surplus = 1/2 * (Difference between minimum price at which he is offering the good and market price * Quantity)

Min price = 1, Market price = 4.5 , Quantity = 5

= 1/2 * (4.5 - 1) * 5

Value of producer surplus = 8.75

3). Total Surplus = Consumer surplus + Producer surplus

= 11.25 + 8.75 = 20